The reaction 2A → B is first order in A with a rate constant of 2.8 × 10−2 s−1 at 80°C. How long (in seconds) will it take for A to decrease from 0.650 M to 0.130 M?

 

1
Expert's answer
2021-03-08T06:04:38-0500

k×t = ln"\frac{A_0}{A_t}"

2.8×10-2 × t = ln "\frac{0.650}{0.130}"

t = (100/2.8) ln 5

t = (100×1.60)/2.8

t = 57.14sec